# LRR-Bench: Left, Right or Rotate? Vision-Language models Still Struggle With Spatial Understanding Tasks

### v2

@@ III-A Absolute Position Reasoning
III-C Can Synthetic Data Benchmark VLMs?
We argue that the answer is yes. For 3D spatial understanding, humans can effortlessly answer these questions. When a model successfully recognizes the image content (i.e., identifies Minecraft elements), its 3D spatial comprehension aligns with that applied to natural images. To evaluate whether VLMs can recognize Minecraft, we select 50 Minecraft screenshots and 50 natural images, tasking the smallest model in each series to distinguish Minecraft images. As shown in table III, all models achieve 100% accuracy except InternVL-7B (96%). This significant performance gap between human and VLMs highlights deficiencies in 3D spatial understanding.
Regarding synthetic natural images, although constrained by auxiliary models, which could make the images simpler than genuine natural images, the models still struggle to answer the corresponding questions accurately. Thus, we posit that synthetic data remains viable for evaluating VLMs’ spatial reasoning capabilities at the current stage.
III-D Evaluation Metrics
@@ IV Experiment
We evaluate multiple state-of-the-art VLMs, including the GPT-4o [17] series, Qwen-VL2 and 2.5 series [28], InternVL2 series [29], Ovis series [46], Janus [47], and Llava-OV series [48]. Due to resource constraints, we utilize the AWQ-quantized versions, which are officially provided, for models exceeding 40B parameters in our evaluation. We also evaluate some fine-tuned VLMs based on 3D datasets, including the llava-3D, SpaceQwen25, and the upgraded llava-3D version based on the Qwen2.5 model. In our evaluation, we do not employ complex reasoning techniques such as Chain of Thought (CoT). We assess two prompting methods: In the first approach, we instruct the model to directly output “Yes” or “No.” In the second approach, we require the model to provide a reasoning process before giving its final response.

### v3

@@ III-A Absolute Position Reasoning
III-C Rationale for the Taxonomy
We categorize spatial understanding into three primary dimensions: movement, rotation, and depth. Within each category, we decompose the problem into granular sub-tasks to maximize coverage of spatial scenarios. Regarding the reviewer’s concern about intermediate concepts like relative depth and containment, our existing benchmark addresses these through specific task designs:
Relative Depth: This requires a model to recognize objects and determine their sequential order along the depth axis. Our Position Sequence task explicitly evaluates this capability. In this task, models must identify not just single images, but the sequential order of images. The cognitive capability required to process image sequences is analogous to the reasoning required for understanding the “front-to-back” ordering of objects in relative depth tasks.
Containment: This involves understanding the inclusion relationship between one object and another. Our Object Movement task addresses this by requiring the model to identify changes in the containment and relative positional relationship between the foreground object and the background environment. This effectively evaluates the core capabilities required for containment understanding.
Overall, our approach isolates specific spatial understanding capabilities through task decomposition. By modeling physical motion as the translation and rotation of rigid bodies (applied to both the object and the observer) and separating foreground from background, our combined tasks provide a comprehensive evaluation of spatial ability.
III-D Alignment with Real-World Perception
We argue that our synthetic benchmark effectively proxies real-world spatial perception for two reasons:
1.
Visual Realism in Static Tasks: For tasks involving Depth and Position, we utilize Flux to generate samples. As demonstrated in fig. 6, these generated images possess a high degree of visual realism comparable to real-world photography.
2.
Geometric Consistency in 3D Tasks: For Movement and Rotation tasks, we design the samples based on the kinematic principles of the real world (i.e., translation and rotation of objects and observers). While the visual domain is synthetic (Minecraft), the spatial logic is identical to reality. To verify that the domain gap (visual recognition) is not the primary cause of failure, we included an “IsMC” control task in the last column of table III. The results show that models achieve near-perfect accuracy in identifying Minecraft objects. Since models can successfully recognize the objects and the spatial relationships follow real-world physics, we conclude that the performance on this benchmark is a valid indicator of real-world spatial reasoning capabilities.
III-E Evaluation Metrics
@@ IV Experiment
We evaluate multiple state-of-the-art VLMs, including the GPT-4o [17] series, Qwen-VL2 and 2.5 series [28], InternVL2 series [29], Ovis series [46], Janus [47], and Llava-OV series [48]. Due to resource constraints, we utilize the AWQ-quantized versions, which are officially provided, for models exceeding 40B parameters in our evaluation. We also evaluate some fine-tuned VLMs based on 3D datasets, including the llava-3D, SpaceQwen25, and the upgraded llava-3D version based on the Qwen2.5 model. In our evaluation, we do not employ complex reasoning techniques such as Chain of Thought (CoT). However, we evaluated two distinct prompting strategies—a direct inquiry and a reasoning-based prompt. These results are presented in table III and table II, respectively. We believe these prompting comparisons serve as a diagnostic intervention, reflecting how users typically interact with these models in practical applications.
@@ IV-D No Single Model or Method Is Universally Superior
IV-E Analysis of Model Failures
Based on the results in table II and table III, we attribute the near-zero scores and general failures to three primary causes:
1.
Insufficient Granularity in Object Recognition: This is particularly evident in smaller models. Analyzing the Pos. and Depth data—the simplest spatial tasks requiring only binary relational identification—some models (e.g., Llava-3D 7B) perform poorly. These models achieve poor performance in other tasks too. Besides, in the Movement tasks, we provided a “Clear” version (denoted by (C)), where distractions were removed to leave only the essential objects. Most models showed improved performance on the Clear version, suggesting the lack of fine-grained object recognition.
2.
Deficiencies in Logical Reasoning: This is observed in the Pos. S (Sequence) and Pos. C (Combination) tasks. To succeed, models must build upon basic positioning capabilities by further understanding the sequential or compositional relationships between sub-images. The significant performance drop in these tasks compared to the basic Pos. task indicates a failure in higher-order logical reasoning.
3.
Incomprehensive Spatial Representations: Models perform poorly on tasks involving intrinsic orientation, such as Obj. H. D. (Heading Direction) and Obj. M. D. (Movement Direction), as well as Camera Rotation. This suggests that current VLMs lack a robust internal representation of 3D geometry, specifically regarding object orientation and the observer’s viewpoint.